OpenStudy (anonymous):

The area of a square is 225 square meters. What is the length of one side of this square? A. 112.5 m B. 15 m C. 50,625 m D. 75 m

OpenStudy (anonymous):

@igotoschool @iambatman

OpenStudy (anonymous):

I think it is c

OpenStudy (anonymous):

Area of a square is \[Area = a^2\] all you have to do is solve for a.

OpenStudy (anonymous):

\[225=a^2\]

OpenStudy (anonymous):

divid it by 2?

OpenStudy (anonymous):

No, you want to get rid of the squared sign, on the a^2 so what would make the exponent on a equal to 1?

OpenStudy (anonymous):

Hint: \[\sqrt{x} \implies x^{1/2}\]

OpenStudy (anonymous):

can u explain little bit better

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok lets say we have a simple problem as such. \[\sqrt{x} = 3\] And I ask you to say solve for x. All you have to do is square both sides: \[\left( \sqrt{x} \right)^2 = 3^2 \implies x^{2/2} = 9 \implies x = 9\]

OpenStudy (alexandervonhumboldt2):